OpenStudy (anonymous):

If 16a^2 x 64=(4 x b)^2 and a and b are positive integers , then b is how many times greater than a?

OpenStudy (anonymous):

Is that "mulitplied by" 64?

OpenStudy (anonymous):

Well, working it out, it must be. What is (4b)(4b) ?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

Okay, I'll just wait till she replies, I have an answer but I'm not sure if it's right.

OpenStudy (anonymous):

yes its multiplied by 64

OpenStudy (anonymous):

Would it be all right to work through the problem? :)

OpenStudy (anonymous):

@leah123 or are you just looking for an answer?

OpenStudy (anonymous):

if you can, can you work out the problem

OpenStudy (anonymous):

Okay :) I'm going to draw this out cuz its easier to understand

OpenStudy (anonymous):

Well, first can you determine what (4b)(4b) is?

OpenStudy (anonymous):

or actually nvm, yes, answer @qpHalcy0n

OpenStudy (anonymous):

thats what the whole thing is equal to

OpenStudy (anonymous):

Darling, you're going to multiply the two together, to get what?

OpenStudy (anonymous):

4 times 4 b times b

OpenStudy (anonymous):

What is 4^2?

OpenStudy (anonymous):

16b^2

OpenStudy (anonymous):

Right, so the 16's go away. The squares go away.

OpenStudy (anonymous):

64a = b. b is how many larger?

OpenStudy (anonymous):

you factor out 16 from everything, if you'd like.

OpenStudy (anonymous):

isn't it 16(a^2 * 4) = 16( b^2)?

OpenStudy (anonymous):

divide by 16 from both sides, and get (a^2 * 4) = (b^2)? and then sqrt both sides? Is that possible @qpHalcy0n

OpenStudy (anonymous):